Find the annual payments to extinguish a debt of P10,000.00 payable for 5years at 12% interest *
1
Expert's answer
2020-11-03T11:28:44-0500

"\bold {Answer}"

Annual Payment, "PMT = \$2,774.10"


"\bold {Solution}"

Assuming the annual payments are made at the end of each year, the question represents an ordinary annuity.

We are a given:

"PV = \$10,000"

"i\% = 12\% \space p.a"

"n = 12 \space years"


The question requires the value of yearly future payments required to clear the $10,000 present debt at 12% annual interest.

For an ordinary annuity,


"PV = PMT \left[\dfrac {1-(1+r)^{-n}}{r} \right]"


"10,000= PMT \left[\dfrac {1-(1+0.12)^{-5}}{0.12} \right]"


"10,000 = PMT \left[\dfrac {1-1.12^{-5}}{0.12} \right]"


"=> PMT = \dfrac {10,000×0.12}{1-1.12^{-5}}"


"= \dfrac {1,200}{0.4325731442}"


"= \$2,774.0973199"


"= \bold {\$2,774.10} \space per \space year"